OpenStudy (anonymous):

Can anyone give me an example of an equation of a line perpendicular and through a point.

OpenStudy (zzr0ck3r):

you mean "Can anyone give me an example of an equation of a line perpendicular to another line and goes through a point."

OpenStudy (anonymous):

I guess so, maybe? the question is written exactly the way i wrote it, and it asking for me to give an example

OpenStudy (zzr0ck3r):

say we have \(y=\frac{1}{2}x+3\) and we want to know the equation of a line that is perpendicular to the line and goes through the point \((0,1)\) the perpendicular part tells us that the line will have slope that is the negative reciprocal of our original functions slope the original functions slope is 1/2 so the slope of our new line will be -2/1 or just -2 so we now have \(y=-2x+b\) so we need to know what b is use the point \((0,1)\) for that let \(x=0\) and \(y=1\) we get \(1=-2x*0+b\implies b=1\) so our line that is perpendicular to \(y=\frac{1}{2}x+3\) through the point \((0,1)\) is \[\large y=-2x+1\]

OpenStudy (zzr0ck3r):

perpendicular by its self means nothing, we must be perpendicular to something... its sort of like saying "to the right of" well to the right of what?

OpenStudy (anonymous):

hmm i see what you mean.

OpenStudy (anonymous):

Thank you, by the way

OpenStudy (zzr0ck3r):

np